### v1

@@ B.1 Image data
In §4.2, we discussed the importance of having a tight $c_{\textrm{lb}}$ so that our measure, $\tilde{\epsilon}$, becomes close to a lower-bound on $\epsilon$-DP, which requires a strong baseline. To strengthen our baseline, we introduce the helper model $h$, which helps the baseline model $b$ by supplying additional features (i.e., embeddings) that can be viewed as side information about the data distribution. The motivation is that $h$’s features might differ between samples from ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$, enhancing the performance of the baseline classifier. This embedding model $h$ is similar in design to $f$ (same task and architecture) but is trained on synthetic data that is close in distribution to the real member data distribution. Whether for the baseline or MIA, we use side information models ($h$ and $f$, respectively) by concatenating the loss of $h(x)$ and $f(x)$ to the final feature representation (more details are provided later) before a last layer of the MIA/Baseline makes the membership the prediction. Since we need labels to compute the loss, we label synthetic images with a Wide ResNet-28-2 in the case of CIFAR10, and a Multi-Label CNN of similar architecture as the target model in the case of CelebA labeling. For both instances, we used a subset of the data, that was used to train the respective generative models, to train the “labeler” classifiers as well.
@@
We use two different modules for each MIA and Baseline training. More precisely, the first module optimizes image classification using a built-in Pytorch ResNet101 classifier. The second module, in the form of a multi-layer perceptron, focuses on classifying member and non-member labels via loss values attributed to these data points by $f$ as input for the loss module of MIA and losses of $e$ to the baseline $b$ respectively. We then stack the scores of both image and loss modules into a logistic regression task (as a form of meta-learning) to get the final outputs for member and non-member data points by MIA and baseline $b$. The MIA and baseline are trained on $4500$ data samples (half members and half generated non-members). The test dataset consists of $10000$ samples, again half members and half generated non-members. The actual and final number of members and non-members that ended up in the test set depends on the Bernoulli samples in our auditing game.
B.2 Language Modeling
@@
Target model. The target model (a small GPT-2 Radford et al. (2019)) task is causal language modeling (CLM), a self-supervised task in which the model predicts the next word in a given sequence of words (i.e., context), which is done on the WikiText-2 dataset Merity et al. (2016), a collection of Wikipedia articles. While the GPT-2 training dataset is undisclosed, Radford et al. (2019) do state that they did not include any Wikipedia document.
@@
One common standard pre-processing in CLM is to break the tokenized training dataset into chunks of equal sizes to avoid padding of the sequences. To achieve this, the entire training dataset is split into token sequences of fixed length, specifically 64 tokens per sequence, refer hereafter as “chunks”. Each of these chunks serves as an individual sample within our training dataset. Therefore, in the experiments, the membership inference will be performed for a chunk, neglecting the weak correlations that could exist between chunks (being from the same article). This chunk-based strategy offers additional benefits in synthetic sample generation. By maintaining uniformity in the length of these synthetic samples, which also consists of 64 tokens, it is possible to effectively mitigate any distinguishability between synthetic and real samples based solely on their length.
To compare our method with the O(1) auditor and another version of our approach using real non-member data, we keep some WikiText samples as non-members, leading the target model not trained on the entire real dataset that we have. Additionally, O(1) method necessitates random inclusion/exclusion of samples into the target model train set Steinke et al. (2023). Hence, we have to include chunks into the $D_{f}$ randomly, which implies that the samples within $D_{f}$ may be from either WikiText-2 or WikiText-103. Although PANORAMIA requires no adjustments to the pipeline, we modify the training pipeline solely for the purpose of comparing it to the O(1) method.
More precisely, we augment $D^{\textrm{te}}_{\textrm{in}}$ by incorporating a few additional documents from WikiText-103, the larger version of the WikiText dataset, into $D_{f}$. We include these additional documents to improve the sample complexity of the hypothesis test in the auditing game, thereby increasing the upper bound of the power of the test. In our experiments with the WikiText dataset, the audit set typically involves 10,000 samples. Assuming 5000 of them are members on average (bearing in mind that our test statistic exclusively considers members), theoretically, we can measure up to 7.21 for either $c_{\textrm{lb}}$ or $\{c\!+\!\epsilon\}_{\textrm{lb}}$.
@@
We collect a total of 72,054 real samples from the WikiText dataset, comprising 36,718 chunks from WT-2 and 35,336 chunks from WT-103. Out of these, 55,282 samples were randomly included in the training set of the target model. We train the target model for 100 epochs in total. Figure 3 displays how the training and validation cross-entropy loss changes throughout the training, when we audit models from epochs 12, 46, and 92. To check how well the target model generalizes, we look at the cross-entropy loss, which is the only metric in a causal language model task to report (or the perplexity Radford et al. (2019) which conveys the same information).
@@
Figure 3: Target model (GPT-2) loss during overtraining on WikiText-2 dataset. We pick GPT-2_E12, GPT-2_E23, GPT-2_E46, GPT-2_E92 as target models to audit corresponding to the gray dashed lines.
@@
Generator. The ${\mathcal{G}}$ is a GPT-2 fine-tuned using a CLM task on dataset $D_{G}$, a subset of $D_{f}$. To create synthetic samples with ${\mathcal{G}}$, we use top-$p$ sampling Holtzman et al. (2019) method in an auto-regressive manner while keeping the generated sequence length fixed at 64 tokens. To make ${\mathcal{G}}$ generate samples less like the real members it learned from (and hopefully, more like real non-members), we introduce more randomness in the generation process in the following manner. First, we set the top_p parameter that determines the probability cutoff for the selection of a pool of tokens to sample from to the value of 1. We also choose a top_k value of 200, controlling the number of available tokens for sampling after applying top_p filtering. In addition, we balance the value of top_k, making it not too small to avoid repetitive generated texts but also not too large to maintain the quality. Finally, we fix the temperature parameter that controls the randomness in the softmax function at one. Indeed, a larger value increases the entropy of the distribution over tokens resulting in a more diverse generated text.
@@
However, the quality of synthetic text depends on the prompts used. Indeed, some prompts, such as "$<|endoftext|>"$, can lead to poor results. To mitigate this issue, we split the $D_{G}$ dataset into two parts: $D_{G}^{train}$ and $D_{G}^{prompt}$. ${\mathcal{G}}$ is fine-tuned on $D_{G}^{train}$, while $D_{G}^{prompt}$ is used for generating prompts. Numerically, the size of the $D_{G}^{train}$ is $44\%$ of the $D_{f}$ and the size of $D_{G}^{prompt}$ is $23\%$ of $D_{f}$. During generation, we sample from $D_{G}^{prompt}$, which is a sequence of length 64. We feed the prompt into the generator and request the generator to generate a suffix of equivalent length. In our experiments, for each prompt, we make 5 synthetic samples, supplying a sufficient number of synthetic non-member samples for different parts of the pipeline of PANORAMIA, including training the helper model, and constructing $D^{\textrm{tr}}_{\textrm{out}}$, $D^{\textrm{te}}_{\textrm{out}}$. Overall, we generate 62524 synthetic samples out of which we use 41968 of them for training the helper model and keep the rest for audit purposes. Note that this generation approach does not inherently favor our audit scheme, and shows how PANORAMIA can leverage existing, public generative models with fine tuning to access high-quality generators at reasonable training cost.
@@
For both the baseline and PANORAMIA classifiers, the training and validation sets consist of 12000 and 2000 samples, respectively, with equal member non-member distribution. The test set consists of 10000 samples, in which the actual number of members and non-members that ended up in the test set depends on the Bernoulli samples in the auditing game. The helper model is fine-tuned for $40$ epochs on synthetic samples and we pick the model that has the lowest validation loss throughout training, generalizing well.
As the GPT-2 component of the classifiers can effectively minimize training loss without achieving strong generalization, regularization is applied to the classifier using weight decay. Additionally, the optimization process is broken into two phases. In the first phase, we exclusively update the parameters associated with the target (in PANORAMIA) or helper model (in the baseline) to ensure that this classifier has the opportunity to focus on these specific features. In the second phase, we optimize the entire model as usual. We train the baseline or MIA with at least 5 seeds, in which the randomness is over the randomness of training (such as initialization of the model, and batch sampling). We select models throughout training with the best validation accuracy. Subsequently, we pick the model with the best validation accuracy among models with different seeds and produce the audit result on the auditing test set with the selected model.
B.3 Tabular data
Target $f$ and the helper model $\mathbf{h.}$ The Target $f$ and the helper $h$ model are both Multi-Layer Perceptron with four hidden layers containing 150, 100, 50, and 10 neurons respectively. Both models are trained using a learning rate of 0.0005 over 100 training epochs. In the generalized case and to obtain the embedding, we retain the model’s parameters that yield the lowest loss on a validation set, which typically occurs at epoch 10. For the overfitted scenario, we keep the model’s state after the 100 training epochs.
@@ B.3 Tabular data
B.4 Comparison with Privacy Auditing with One (1) Training Run: Experimental Details
We implement the black-box auditor version of O(1) approach Steinke et al. (2023). This method assigns a membership score to a sample based on its loss value under the target model. They also subtract the sample’s loss under the initial state (or generally, a randomly initialized model) of the target model, helping to distinguish members from non-members even more. In our instantiation of the O(1) approach, we only consider the loss of samples on the final state of the target model. Moreover, in their audit, they choose not to guess the membership of every sample. This abstention has an advantage over making wrong predictions as it does not increase their baseline. Roughly speaking, their baseline is the total number of correct guesses achieved by employing a randomized response $(\epsilon,0)$ mechanism, for those samples that O(1) auditor opts to predict. We incorporate this abstention approach in our implementation by using two thresholds, $t_{+}$ and $t_{-}$. More precisely, samples with scores below $t_{+}$ are predicted as members, those above $t_{-}$ as non-members, and the rest are abstained from prediction. We check all possible combinations of $t_{+}$ and $t_{-}$ and report the highest $\epsilon$ among them, following a common practice Zanella-Béguelin et al. (2022); Maddock et al. (2023). We also set $\delta$ to 0 and use a confidence interval of 0.05 in their test. In PANORAMIA, for each hypothesis test (whether for $c_{\textrm{lb}}$ or $\{c\!+\!\epsilon\}_{\textrm{lb}}$), we stick to a 0.025 confidence interval for each one, adding up to an overall confidence level of 0.05. Furthermore, the audit set of O(1) is the same as the audit set of PANORAMIA.
B.5 Comparison with MIA on RM; RN: Experimental details on tabular
In order to proceed with the comparison of PANORAMIA with a variant of PANORAMIA in which we use real data instead of synthetic data, we need a significant number of real non-member observations. Due to the limited size of the adult dataset ($\approx 30k$ observations), we train the target model $f$ on only half of the dataset and use the remaining half to train and evaluate the MIA.
Appendix C Results: More experiments and Detailed Discussion
C.1 Baseline Strength Evaluation on Text and Tabular datasets
@@
CelebA Baselineno helper 0.91
@@
WikiText-2 Baselineno helper 2.34
[…]
Adult Baselineno helper 2.01
@@
The basis of our audit directly depends on how well our baseline classifier distinguishes between real member data samples and synthetic non-member data samples generated by our generative model. To assess the strength of the baseline classifier, we train the baseline’s helper model $h$ on two datasets, synthetic data, and real non-members. We also train a baseline without any helper model. This enables us to identify which setup allows the baseline classifier to distinguish the member and non-member data the best in terms of the highest $c_{lb}$ value ($c_{lb}$ is reported for each case in table 6). For initial analysis, we keep a sub-set of real data as non-members to train a helper model with them, for conducting this ablation study on the baseline. Our experiments confirm that the baseline with loss values from a helper model trained on synthetic data performs better compared to a helper model trained on real non-member data. This confirms the strength of our framework in terms of not having a dependency of using real non-members. Therefore, we choose this setting for our baseline helper model. In addition, we also experiment with different helper model architectures in the case of CIFAR10, looking for the best baseline performance among them.
@@
For each data modality setting, we also train the baseline classifier on an increasing dataset size of real member and synthetic non-member data points to gauge the trend of the baseline performance as the training dataset size increases (see Figure 2 and Figure LABEL:fig:baseline-othermodalities-eval). This experiment allows us to aim for a worst-case empirical lower bound on how close the generated non-member data is to the real member data distribution, i.e. what would be the largest $c_{\textrm{lb}}$ we can reach as we increase the training set size. Subsequently, we can compare the $c_{\textrm{lb}}$ reported in table 4 (corresponding to the vertical dashed lines in Figure 2 and Figure LABEL:fig:baseline-othermodalities-eval) to the largest $c_{\textrm{lb}}$ we can achieve in this experiment. In the following, we explain why we choose that number of training samples for our baseline in the main results.
On the WikiText dataset, we augment the baseline’s training set by shortening the validation and test sets (from 1000 to 500 and from 10000 to 2000, respectively). This modification accounts for the mismatch between the $c_{\textrm{lb}}$ at the vertical dashed line in Figure LABEL:subfig:nlp_plateau and the reported one in Table 4. Nevertheless, the maximum average $c_{\textrm{lb}}$ attained on our largest training set for the baseline (represented by the horizontal dashed line in Figure LABEL:subfig:nlp_plateau) is 3.66, which is still lower than 3.78, the one reported in Table 4.
Due to dataset size constraints in practicality, in terms of how much real member data is available, the train dataset size for our MIA and baseline where the performance of the baseline in terms of $c_{lb}$ is the highest. This accounts for how much data we have in terms of member data for the target model, and the split we account for while training the respective generative models, as well as the real members test data used to evaluate the MIA and baseline. These data constraints lead us to choose the train dataset size as shown by the vertical line in Figure 2 and Figure LABEL:fig:baseline-othermodalities-eval. Hence the final values reported in Table 3 are on these corresponding data sizes. It is also important to note that the MIA is expected to show similar behavior if we use a larger dataset size for training if it was available.
The experiments are repeated ten times to account for randomness during training (e.g., initialization, sampling in batches) and report the results over a $95\%$ confidence interval.
C.2 Further Analysis on General Auditing Results

### v2

@@ C.1 Image data
In §4.2, we discussed the importance of having a tight $c_{\textrm{lb}}$ so that our measure, $\tilde{\epsilon}$, becomes close to a lower-bound on $\epsilon$-DP, which requires a strong baseline. To strengthen our baseline, we introduce the helper model $h$, which helps the baseline model $b$ by supplying additional features (i.e., embeddings) that can be viewed as side information about the data distribution. The motivation is that $h$’s features might differ between samples from ${\mathcal{D}}$ and ${\mathcal{D}}^{\prime}$, enhancing the performance of the baseline classifier. This embedding model $h$ is similar in design to $f$ (same task and architecture) but is trained on synthetic data that is close in distribution to the real member data distribution.
The helper model has the same classification task and architecture as the target model. To train it, we generate separate sets of training and validation data using our generator. For image data, the synthetic samples do not have a label. We thus train a labeler model, also a classifier with the same task, on the same dataset used to train the generator. We use the labeler model to provide labels for our synthetic samples (training and validation sets above). We train the helper model on the resulting training set, and select hyperparameters on the validation set.
Whether for the baseline or MIA, we use side information models ($h$ and $f$, respectively) by concatenating the loss of $h(x)$ and $f(x)$ to the final feature representation (more details are provided later) before a last layer of the MIA/Baseline makes the membership the prediction. Since we need labels to compute the loss, we label synthetic images with a Wide ResNet-28-2 in the case of CIFAR10, and a Multi-Label CNN of similar architecture as the target model in the case of CelebA labeling. For both instances, we used a subset of the data, that was used to train the respective generative models, to train the “labeler” classifiers as well. The labeler used to train the helper model for image data has 80.4% test accuracy on the CIFAR10 real data test set. The rationale for this approach is to augment the baseline with a model providing good features (here for generated data) to balance the good features provided to the MIA by $f$ outside of the membership information. In practice, the labeled generated data seems enough to provide such good features, despite the fact that the labeler is not extremely accurate. We studied alternative designs (e.g., a model trained on non-member data, no helper model) in Appendix D.1, Table 5, and the helper model trained on the synthetic data task performs best (while not requiring non-member data, which is a key point).
@@
We use two different modules for each MIA and Baseline training. More precisely, the first module optimizes image classification using a built-in Pytorch ResNet101 classifier. The second module, in the form of a multi-layer perceptron, focuses on classifying member and non-member labels via loss values attributed to these data points by $f$ as input for the loss module of MIA and losses of $e$ to the baseline $b$ respectively. We then stack the scores of both image and loss modules into a logistic regression task (as a form of meta-learning) to get the final outputs for member and non-member data points by MIA and baseline $b$. When training the baseline and MIA models, we use a validation set to select hyper-parameters and training stopping time that maximize the lower bounds (effectively maximizing the $c_{lb}$ or $\tilde{\epsilon}_{lb}$ for the Baseline and MIA respectively) on the validation set. The MIA and baseline are trained on $4500$ data samples (half members and half generated non-members). The test dataset consists of $10000$ samples, again half members and half generated non-members. The actual and final number of members and non-members that ended up in the test set depends on the Bernoulli samples in our auditing game. We repeat the training process over 5 times, each time independently resampling the training dataset for the MIA and baseline (keeping the train dataset size fixed). We report all our results over a $95\%$ confidence interval and report the standard deviation for our results in Table 3. The compute resources used for MIA and Baseline training were mainly running all experiments on cloud-hosted VMs, using 1 v100l GPU, 4 nodes per task, and 32G memory requested for each job on the cloud cluster. The time to run the MIA and baseline attack pipeline was around 10 hours including hyperparameter tuning on $\epsilon_{lb}$ and $c_{lb}$ respectively.
C.2 Language Modeling
@@
Target model. The target model (a small GPT-2 Radford et al. (2019)) task is causal language modeling (CLM), a self-supervised task in which the model predicts the next word in a given sequence of words (i.e., context), which is done on a subset of WikiText-103 dataset Merity et al. (2016), a collection of Wikipedia articles. While the GPT-2 training dataset is undisclosed, Radford et al. (2019) do state that they did not include any Wikipedia document.
@@
One common standard pre-processing in CLM is to break the tokenized training dataset into chunks of equal sizes to avoid padding of the sequences. To achieve this, the entire training dataset is split into token sequences of fixed length, specifically 64 tokens per sequence, refer hereafter as “chunks”. Each of these chunks serves as an individual sample within our training dataset. Therefore, in the experiments, the membership inference will be performed for a chunk. In other words, we adopt an example-level differential privacy (DP) viewpoint, where an example is considered a sequence of tokens that constitute a row in a batch of samples. This chunk-based strategy offers additional benefits in synthetic sample generation. By maintaining uniformity in the length of these synthetic samples, which also consists of 64 tokens, it is possible to effectively mitigate any distinguishability between synthetic and real samples based solely on their length. We also consider the possibility of weak correlations between chunks (e.g., being from the same article). To mitigate this, we perform a form of stratified sampling as a pre-processing step. From each document, we include only the first 50 chunks and discard the rest. We select 2,113 documents in total, resulting in 105,650 chunks in our dataset.
We split the dataset into train and test with a 90:10 ratio for the target model (We insist on following the typical training pipelines in machine learning, since we plan to conduct post-hoc audits with PANORAMIA). The test dataset provides real non-member samples, allowing comparison to the O(1) auditor and another version of our approach which uses real non-member data. The training dataset ($D_{f}$) contains 95,085 samples, while the test dataset consists of 10,565 samples.
@@
We train the target model for 200 epochs in total. Figure 13 displays how the training and validation cross-entropy loss changes throughout the training, when we audit models from epochs 37 (best generalization), 75, and 150. To check how well the target model generalizes, we look at the cross-entropy loss (on a validation set), which is the only metric in a causal language model task to report (or the perplexity Radford et al. (2019) which conveys the same information).
@@
Figure 13: Target model (GPT-2) loss during overtraining on WikiText dataset. We pick GPT-2_E37, GPT-2_E75, GPT-2_E150 as target models to audit corresponding to the gray dashed lines.
The compute resources used for training the target models are 4 V100-32gb GPUs, running on cloud-hosted VMs. Using this setup, fine-tuning the target model for 200 epochs required about 13 hours.
@@
Generator. The ${\mathcal{G}}$ is a GPT-2 fine-tuned using a CLM task on dataset $D_{G}$, a subset of $D_{f}$. To create synthetic samples with ${\mathcal{G}}$, we use top-$k$ sampling Holtzman et al. (2019) method in an auto-regressive manner while keeping the generated sequence length fixed at 64 tokens. To make ${\mathcal{G}}$ generate samples less like the real members it learned from (and hopefully, more like real non-members), we introduce more randomness in the generation process in the following manner. We choose a top_k value of 200, controlling the number of available tokens for sampling. In addition, we balance the value of top_k, making it not too small to avoid repetitive generated texts but also not too large to maintain the quality. Finally, we fix the temperature parameter that controls the randomness in the softmax function at one. Indeed, a larger value increases the entropy of the distribution over tokens resulting in a more diverse generated text.
@@
However, the quality of synthetic text depends on the prompts used. To mitigate this issue, we split the $D_{G}$ dataset into two parts: $D_{G}^{train}$ and $D_{G}^{prompt}$. ${\mathcal{G}}$ is fine-tuned on $D_{G}^{train}$, while $D_{G}^{prompt}$ is used for generating prompts. Numerically, the size of the $D_{G}^{train}$ is $35\%$ of the $D_{f}$ and the size of $D_{G}^{prompt}$ is $15\%$ of $D_{f}$. During generation, we sample from $D_{G}^{prompt}$, which is a sequence of length 64. We feed the prompt into the generator and request the generator to generate a suffix of equivalent length. In our experiments, for each prompt, we make 8 synthetic samples, supplying a sufficient number of synthetic non-member samples for different parts of the pipeline of PANORAMIA, including training the helper model, and constructing $D^{\textrm{tr}}_{\textrm{out}}$, $D^{\textrm{te}}_{\textrm{out}}$. Overall, we generate 92276 synthetic samples out of which we use 53502 of them for training the helper model and keep the rest for audit purposes. Note that this generation approach does not inherently favor our audit scheme, and shows how PANORAMIA can leverage existing, public generative models with fine tuning to access high-quality generators at reasonable training cost.
We also perform a sanity check on the quality of the generated data. We visualize the loss distribution of in-distribution members, in-distribution non-members, and synthetic data under both the target and helper models, aiming for synthetic data that behaves similarly to in-distribution non-member data. LABEL:fig:gpt_losses visualizes these distributions for our target and helper models.
The compute resources used for fine-tuning the generator model, and synthetic text generation are 4 V100-32gb GPUs, running on cloud-hosted VMs. Using this setup, fine-tuning the GPT2-small model took approximately 1.5 hours, and generating the synthetic samples required about 1 hour.
@@
In the main results in Table 3, for both the baseline and PANORAMIA classifiers, the training and validation sets consist of 20000 and 2000 samples, respectively, with equal member non-member distribution. The test set consists of 10000 samples, in which the actual number of members and non-members that ended up in the test set depends on the Bernoulli samples in the auditing game. The helper model is fine-tuned for $60$ epochs on synthetic samples and we pick the model that has the lowest validation loss throughout training, generalizing well.
As the GPT-2 component of the classifiers can effectively minimize training loss without achieving strong generalization, regularization is applied to the classifier using weight decay. Additionally, the optimization process is broken into two phases. In the first phase, we exclusively update the parameters associated with the target (in PANORAMIA) or helper model (in the baseline) to ensure that this classifier has the opportunity to focus on these specific features. In the second phase, we optimize the entire model as usual. We train both the baseline and Membership Inference Attack (MIA) models with 5 different seeds. The randomness is over: 1. The data split into train/validation/test sets for both the baseline and MIA classifiers. Each classifier is retrained for every seed. 2. The coin flips in our auditing game (see Definition 2). For a fixed seed mentioned above, we employ an additional set of 5 seeds to account for the randomness in the training algorithm (such as initialization of the model, and batch sampling). We then train a classifier for each of these seeds. From the 5 models generated for each latter seed, we select the one with the largest $c_{\textrm{lb}}$ value (for the baseline) or $\{c\!+\!\epsilon\}_{\textrm{lb}}$ (for MIA), determined over the validation set. We calculate the 95% confidence interval of our measurements using the t-score.
The compute resources used for MIA and Baseline training were 1 GPU with 32gb memory, running on cloud-hosted VMs. The compute time depends on the training set size, for the largest training size, it took about 3 hours to train the baseline or MIA.
The compute resources used for fine-tuning the helper model are 4 V100-32gb GPUs, running on cloud-hosted VMs. Using this setup, it took about 4 hours to fine-tune the helper model.
C.3 Tabular data
Target $f$ and the helper model $\mathbf{h.}$ The Target $f$ and the helper $h$ model are both Multi-Layer Perceptron with four hidden layers containing 150, 100, 50, and 10 neurons respectively. Both models are trained using a learning rate of 0.0005 over 100 training epochs. In the generalized case and to obtain the embedding, we retain the model’s parameters that yield the lowest loss on a validation set, which typically occurs at epoch 10 (MLPE_10). For the overfitted scenario, we keep the model’s state after the 100 training epochs (MLP_E100). The non overfitted model and the overfitted one achieved an accuracy of 86% and 82% respectively.
@@ C.3 Tabular data
C.4 Comparison with Privacy Auditing with One (1) Training Run: Experimental Details
We implement the black-box auditor version of O(1) approach Steinke et al. (2023). This method assigns a membership score to a sample based on its loss value under the target model. They also subtract the sample’s loss under the initial state (or generally, a randomly initialized model) of the target model, helping to distinguish members from non-members even more. In our instantiation of the O(1) approach, we only consider the loss of samples on the final state of the target model. Moreover, in their audit, they choose not to guess the membership of every sample. This abstention has an advantage over making wrong predictions as it does not increase their baseline. Roughly speaking, their baseline is the total number of correct guesses achieved by employing a randomized response $(\epsilon,0)$ mechanism, for those samples that O(1) auditor opts to predict. We incorporate this abstention approach in our implementation by using two thresholds, $t_{+}$ and $t_{-}$. More precisely, samples with scores below $t_{+}$ are predicted as members, those above $t_{-}$ as non-members, and the rest are abstained from prediction. We check for different combinations of $t_{+}$ and $t_{-}$ and report the highest $\epsilon$ among them. This involves performing multiple tests on the same evidence, each with a separate confidence level of 95%. To ensure that these tests hold collectively, we apply a union bound and adjust the significance level accordingly. The total number of separate tests depends on the different number of guesses that O(1) can make. We also set $\delta$ to 0 and use a confidence interval of 0.05 in their test. In PANORAMIA, for each hypothesis test (whether for $c_{\textrm{lb}}$ or $\{c\!+\!\epsilon\}_{\textrm{lb}}$), we stick to a 0.025 confidence interval for each one, adding up to an overall confidence level of 0.05.
Appendix D Results: Additional Experiments and Detailed Discussion
D.1 Baseline Strength Evaluation on Text and Tabular datasets
@@
CIFAR-10 Baselineno helper $1.25\pm 0.24$
@@
WikiText Baselineno helper $3.11\pm 0.15$
@@
The basis of our privacy measurement directly depends on how well our baseline classifier distinguishes between real member data samples and synthetic non-member data samples generated by our generative model. As mentioned in Section 5.1 and Section C.1, in order to increase the performance of our baseline $b$, we mimic the role of the target model $f$’s loss in the MIA using a helper model $h$, which adds a loss-based feature to $b$. To assess the strength of the baseline classifier, we train the baseline’s helper model $h$ on two datasets, synthetic data, and real non-members. We also train a baseline without any helper model. This enables us to identify which setup allows the baseline classifier to distinguish the member and non-member data the best in terms of the highest $c_{lb}$ value ($c_{lb}$ is reported for each case in table 5). For initial analysis, we keep a sub-set of real data as non-members to train a helper model with them, for conducting this ablation study on the baseline. Our experiments confirm that the baseline with loss values from a helper model trained on synthetic data performs better compared to a helper model trained on real non-member data. This confirms the strength of our framework in terms of not having a dependency of using real non-members. Therefore, we choose this setting for our baseline helper model. In addition, we also experiment with different helper model architectures in the case of CIFAR10, looking for the best baseline performance among them. For the WikiText dataset, the $c_{\textrm{lb}}$ reported in Table 5 differs from that in Table 3 because we had less data available for this experiment. We withheld half of the real samples to train the helper model with real samples in its training set.
@@
For each data modality setting, we also train the baseline classifier on an increasing dataset size of real member and synthetic non-member data points to gauge the trend of the baseline performance as the training dataset size increases (see Figure 2 and Figure LABEL:fig:baseline-other-modalities-eval). This experiment allows us to aim for a worst-case empirical lower bound on how close the generated non-member data is to the real member data distribution, i.e. what would be the largest $c_{\textrm{lb}}$ we can reach as we increase the training set size. Subsequently, we can compare the $c_{\textrm{lb}}$ reported in table 3 (corresponding to the vertical dashed lines in Figure 2 and Figure LABEL:fig:baseline-other-modalities-eval) to the largest $c_{\textrm{lb}}$ we can achieve in this experiment. In the following, we explain why we chose that number of training samples for our baseline in the main results.
On the WikiText dataset, we increase the baseline’s training set size up to 50k, while fixing the validation and test sets sizes to 2000 and 10000, respectively). LABEL:fig:baseline-other-modalities-eval shows the changes of $c_{\textrm{lb}}$ while varying the training set size. The vertical dashed line specified the number of training samples we chose for our main auditing results reported in Table 3.
Now, let us discuss why we don’t choose the largest training set size available for the baseline. The measurement for $c_{\textrm{lb}}$ depends not only on the training set size but also on the test set size. Therefore, we need to balance the sizes of the train and test sets. We conducted another experiment where we varied the train set size and, instead of fixing the test set size, allocated the remaining available data to the test set (keeping in mind that our data for the audit is limited by the number of real members allocated for audit purposes).LABEL:subfig:nlp_vary_train_test depicts the trend of changes. As expected, we observed an initial increase in $c_{\textrm{lb}}$ as the training set size increased (and the number of test data is sufficient to capture that). However, it decreased toward the end because of the smaller test set size. For a train set size of 40k, we found the best trade-off between train and test set sizes, which justifies our choice of 40k for the training set size in the main results. The same argument applies for other data modalities.
Let us make one last point. It is also important to note that the MIA is expected to show similar behavior if we use a larger train set size or test set size. Hence, the gap between the performance of MIA and baseline (which is our privacy measurement) is not expected to be affected (see Section D.4).
The experiments are repeated five times, and we provide a 95% confidence interval. For the image data modality, the randomness lies in the training data provided to the classifiers, while the test set remains fixed throughout. For the text data modality, both the train and test sets change across the five runs (the test set size is fixed to 10k).
D.2 Further Analysis on General Privacy Measurement Results